Use separation of variables to solve the initial value problem. 3x2 and y = 1 when x = 0 21) y =
Answer #1

2 2 da Y= 3x sy Study = 3 yl 3½ 3x 3 = tc 2 y 3/2 = x²tc y312 = 3 x + c Now x=0, y=1 (13/2 Fotc c=1 sol 13/2 x+1 Answes